OpenStudy (anonymous):

can someone check this for me? Square root 6a - 4square root 54a - 4 square root 216a simplify by combining like radical terms. Assume all veriables and radicands represent nonnegative numbers

OpenStudy (anonymous):

sqrt(6a) -4*sqrt(54a) -4*sqrt(216a) 54=9*6 & 9=3^2, so sqrt(54)=3*sqr(6) 216=36*6 & 36=6^2, so sqrt(216)=6*sqrt(6) yu hav sqrt(6a)-12*sqrt(6a) -24*sqrt(6a) -35*sqrt(61)

jhonyy9 (jhonyy9):

@glea finaly is -35sqrt6a and not sqrt61 ok ? ATTENTION !!!